• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 03:27:21 on July26,2005

    Entry number 149484

    keyword=MCC update

    Ken gave us another update.  There was a problem with a PSS IOC which has to be 
    fixed (and perhaps is already fixed) and then they can try to finish up the 
    pathlength work.  If the pathlength work doesn't go well after another hour, 
    they will deliver beam as it was before the work started and try to get things 
    fixed during the day shift.  If the Compton tune is still bad, we'll take data 
    but NOT run the Comptom.